The largest angle of a triangle is 111 degrees greater than the smallest angle. The sum of the two smaller angles is 68 degrees less than the largest angle. Find the measure of each angle.
Answer by Paul(988)

111+x+x+y=180
2x+y=69
y=69-2x
x+69-2x=111+x-68 (simplify)
-2x=-26
x=13
111+13=124
13+y=124-68
y=43
Hence, the 3 angles in the triangle are 13, 43, and 124 degrees.